What volume (in milliliters) of a 0.315 M NaOH solution contains 6.22 g of NaOH?
1
Expert's answer
2018-01-22T05:58:34-0500
m(NaOH) = 6.22g
Cm = 0.315M
n(NaOH) = m(NaOH)/M(NaOH) = 6.22g/40 g/mol = 0.156g
Cm = n(NaOH)/V(NaOH)
V(NaOH) = n(NaOH)/ Cm = 0.156g/0.315 M = 0.49 L